OpenStudy (anonymous):

Geometry Help?! An oblique cone has a height of 5 m ad a base with a diameter of 18 m. A right cone has a 5 m height and a radius of 9 m for its base. Which describes the volume of the oblique cone (Vo) compared with the volume of the right cone (Vr)? A.) Vr

OpenStudy (yttrium):

For this problem, you can see that the height of the two cones are equal. Also, the radius of the oblique cone is twice of that the right cone.

OpenStudy (yttrium):

From the general formula for volume of the cone. We have\[V = \frac{ 1 }{ 3 } \pi r^{2} h\]

OpenStudy (anonymous):

okay

OpenStudy (yttrium):

Okay. From the facts I said on my first comment. We can say now that \[h_{o} = h _{r}\] and \[r _{o} = 2r _{r}\]

OpenStudy (yttrium):

We can determine the relation of their volume by using proportion. That is by solving Vo/Vr

OpenStudy (yttrium):

From that, \[\frac{ V _{o} }{ V _{r} } = \frac{ \frac{ 1 }{ 3 } \pi (2r _{r})^{2}h _{o} }{ \frac{ 1 }{ 3 } \pi (r _{r})^{2} h _{o}}\]

OpenStudy (yttrium):

Simplifying the above equation, we can see that \[\frac{ Vo }{ Vr } = 2\]